Let X∼Bin(6,0.7) and set Y=(X−2)^2 . Find P(Y=1).
Answer #1

P(Y = 1)

= P((X - 2)^2 = 1)

= P(X-2 = 1 or -1)

= P(X = 3 or X = 1)

= P(X = 1) + P(X = 3)

{\displaystyle f(k,n,p)=\Pr(k;n,p)=\Pr(X=k)={\binom {n}{k}}p^{k}(1-p)^{n-k}}

n = 6 , p = 0.7

P(X = 1) = 6C1 0.7 * 0.3^5

=6 * 0.7 * 0.3^5

= 0.010206

P(X = 3) = 6C3 * 0.7^3 * 0.3^3

= 20* 0.7^3 * 0.3^3

= 0.18522

hence required probability

= 0.18522 + 0.010206

= 0.195426

Please rate